Circular No. 8817

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2004 (45) 1 F. Marchis and M. Baek, University of California at Berkeley; and P.

Descamps, J. Berthier, D. Hestroffer, and F. Vachier, Institut de Mecanique Celeste et de Calcul des Ephemerides, Paris, report the discovery of a new satellite orbiting (45) Eugenia. S/2004 (45) 1 was detected after careful analysis of three observations collected with K-band filters using the Very Large Telescope ‘YEPUN’ and its adaptive-optics system (NACO) on 2004 Feb. 14.15404, 15.14620, and 16.15435 UT, the satellite appearing at a distance of ∼ 0′′.4 from the primary in p.a. 156^{o}, 321^{o}, and 124^{o}, respectively. Based on the integrated brightness ratio of ∼ 7.9 between the satellite and the primary, the diameter of S/2004 (45) 1 is estimated to be ∼ 6 km. The orbital parameters of this satellite give a mass for the system in agreement with the mass derived from the study of previously known satellite, (45) Eugenia I (Petit-Prince; cf. IAUCs 7129, 7503).

COMET 186P/2007 B3 (GARRADD) Further to IAUC 8801 and 8806, M. Meyer (Limburg, Germany) re-

ports that he has identified this comet on Digitized Sky Survey plates taken with the U.K. Schmidt telescope on 1975 May 31 and June 3 (two red plates each; centers of trails measured) and on 1996 Feb. 26 (single red plate; comet trail involved with star, ends of trail measured). In verifying these identifications, G. V. Williams (Minor Planet Center) then identified P/2007 B3 with an 18th-magnitude comet reported by R. D. Eberst (cf. IAUC 3247) one year after the U.K. Schmidt plates of 1977 July 18 and 19 were taken (later designated X/1977 O1 in the Catalogue of Cometary Orbits 1995). Meyer’s astrometry and the precessed Eberst astrometry appear on MPEC 2007-D71 with new orbital elements and residuals; these observations at multiple apparitions permit the numbering of this comet as 186P/Garradd (cf. MPC 58997).

SUPERNOVA 2007af IN NGC 5584 S. Nakano, Sumoto, Japan, reports the discovery by K. Itagaki (cf.

IAUC 8792) of a fairly bright supernova located at α = 14^{h}22^{m}21^{s}.03, δ = −0^{o}23′37′′.6 (equinox 2000.0), which is 40′′ west and 22′′ south of the center of NGC 5584. Unfiltered CCD magnitudes for 2007af from Itagaki: Feb. 24.81 UT, [19.0; 25.73, [18.5; Mar. 1.84, 15.4 (discovery); 2.576, 15.3; 7.648, 13.7. SN 2007af is a type-ia supernova that was discovered well before maximum light (spectroscopic details on CBET 865).

2007 March 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green